investments Annie and michael invested $8000 last year in tax-free bonds. some of the bonds earned 8% simple interest while the rest earned 6% simple interest. at the end of the year they had earned $580 in interest. how much did they invest at each interest rate?
Answer by rfer(10425)

0.08x+0,06(8000-x)=580
0.08x+480-0.06x=580
0.02=100
x=$5000 at8%
8000-x=$3000 at 6%